OpenStudy (anonymous):

The sides of a triangular lot are represented by x, 3x, and 3x + 2. Find each side if the perimeter is 632 feet. The lengths of the sides of the triangle are ,___ ,__ and .

OpenStudy (kropot72):

x + 3x + (3x + 2) = 632 7x + 2 = 632 7x = 632 - 2 = 630 Can you solve for x now?

OpenStudy (kropot72):

@Claire4christ Are you there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

632-2??

OpenStudy (kropot72):

7x = 630 x = 630/7 = ? feet

OpenStudy (anonymous):

90

OpenStudy (kropot72):

Correct! So the other two sides are: \[3\times 90=?feet\] \[(3\times 90)+2=?feet\]

OpenStudy (anonymous):

272

OpenStudy (kropot72):

Yes. 270 feet is the length of another side. How about the third side which is \[3\times 90=?feet\]

OpenStudy (anonymous):

90,270,272

OpenStudy (kropot72):

Good work! Don't forget to put the units (feet) after each length. 90 feet, 270 feet and 272 feet.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

I got another question??

OpenStudy (kropot72):

You're welcome :)

OpenStudy (kropot72):

If you post another question I'll take a look.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do you post another question.

OpenStudy (kropot72):

You need to close this one before you can post another question.